# with theoretical and observational applications

### v1

@@ Abstract
We focus on minimally coupled (multi)field quintessence models and their realistic solutions. In a model-independent manner, we describe analytically these cosmological solutions throughout the universe history. Starting with a kination - radiation domination phase, we obtain an upper bound on the scalar potential to guarantee an early kination: $V(\varphi)\ll e^{-\sqrt{6}\varphi}$. Turning to the radiation - matter phase, we obtain analytical expressions for the scale factor $a(t)$ (not $t(a)$) and the scalar fields $\varphi^{i}(t)$ (usually neglected). These allow us to evaluate analytically the freezing of scalar fields, typically $\Delta\varphi\lesssim 10^{-2}$, as well as the transition moment of the dark energy equation of state parameter $w_{\varphi}$ from $+1$ to $-1$, with excellent agreement to the numerics. We comment on this freezing in view of string theory model building, and of some cosmological events. Turning to the latest phase of matter - dark energy domination, we show that the (multi)field displacement is sub-Planckian: $\Delta\varphi\leq 1$. We also provide for that phase analytic expressions for $\int(w_{\varphi}+1)\,{\rm d}N$; we relate those to observational targets that we propose. Turning to the CPL parametrisation, while discussing a phantom behaviour, we derive analytical bounds on $w_{0}$ and $w_{a}$.
@@ 1 Introduction and results summary
For about a century, we have known that our universe is expanding. Less than thirty years ago, the present expansion was observed to be accelerating. This acceleration is due to a so-called dark energy, which behaves very differently than ordinary matter or radiation. What is the nature of dark energy? This open problem is a key question of modern fundamental physics. A standard proposal is to describe dark energy as a constant energy density given by a cosmological constant $\Lambda>0$. This has led to a concordance cosmological model named $\Lambda$CDM, well-tested and in fairly good agreement with observations. Another option however is to have an energy density varying with time, giving a dynamical dark energy. One realisation of this possibility is through so-called quintessence models [1, 2, 3, 4] (see e.g. [5, 6, 7] for early reviews and constraints), where the dynamics is that of scalar fields evolving in a scalar potential. Recent (and coming) observations, as well as theoretical arguments, have revived the possibility of explaining dark energy by quintessence. In this work, we perform a model-independent analysis of realistic quintessence scenarios, identifying interesting theoretical features for model building, as well as characteristic signatures that could possibly be observed, helping to distinguish such models from $\Lambda$CDM.
On the observational side, the first results of the DESI collaboration have recently been released [8]. Those are based on collected data until redshift $z=4$ with an unprecedent precision level. Consistently with the recent DES results [9], the DESI observational results are compatible with a dynamical dark energy model. More DESI results are announced for the coming years, and further experiments (Euclid, LSST/Vera Rubin) will provide independent measurements to similar precision. On the theoretical side, difficulties have been pointed-out in obtaining a universe with a positive cosmological constant (a.k.a. a de Sitter spacetime) from string theory, a quantum gravity theory candidate to be a fundamental theory of Nature. These difficulties have been characterised in terms of the strong de Sitter conjecture of the swampland program [10, 11]. While not claimed to be impossible, there is no known well-controlled string theory construction of such a solution up-to-date. As an alternative, there has been a recent interest in rather realising quintessence models from string theory (see [12, 13, 14, 15, 16, 17, 18, 19] for a recent sample, and [15] for further references). While necessary ingredients are naturally present in string effective theories, finding the right dynamics and an appropriate scalar potential is not an easy task. But a lot of activity is currently dedicated to these matters, including cosmological model building and related investigations [20, 21, 22, 23, 24, 25].
@@
indicating a dominant dark energy component, that provides acceleration. We consider in Section 3 three different models, and for each of them we obtain a realistic solution. The first one is $\Lambda$CDM, which can be viewed as a particular case of (1.1), and the other two are single field quintessence models, with respectively an exponential or a hilltop scalar potential. For quintessence models, obtaining a realistic solution (especially with a radiation domination phase) requires a tuning of the initial conditions, which we achieve. The evolution of the $\Omega_{n}$ is then similar from one model to the other, and we display it in Figure 1. We indicate in Figure 1(a) that different phases in this universe history can be delimited by moments, expressed in terms of $N$, that characterise equality ($N_{{\rm kin}r},N_{rm},N_{m\varphi}$) or maxima ($N_{r},N_{m}$) of the $\Omega_{n}$. We discuss these phases and specific moments in great details in Section 3.4, and provide various related analytic expressions. Given some values today $\Omega_{n0}$, it can be seen in Figure 1(b) that some of these moments are necessarily different between $\Lambda$CDM and quintessence, leading us to introduce extra notations: for instance, we label the moment for the maximum of $\Omega_{m}$ as $N_{m\,\Lambda}$ or $N_{m\,q}$ respectively. Similarly for the equality moment, we distinguish $N_{m\varphi\,\Lambda}$ and $N_{m\varphi\,q}$.
@@
The evolution, depicted in Figure 2(b), indicates the initial domination of kinetic energy ($w_{\varphi}=1$) as in kination, followed by a domination of the scalar potential ($w_{\varphi}=-1$) as for a cosmological constant, and a final rise towards the value today. This recent evolution is the prime target of DESI observations and will be of interest here; in addition the transition moment at which $w_{\varphi}=0$ will be determined analytically. Note that the latest part of this evolution is that of thawing quintessence models, which are included in this study.
@@
The analytical solutions obtained provide a better understanding of the physics of the scalar fields; that of the scale factor is nothing but a continuous growth at various rates. Provided some initial kinetic energy and a negligible potential, the fields start by rolling up or down the potential: this is kination. The Hubble friction being huge in the early universe, the fields are quickly slowed down. During radiation domination, the kinetic energy continues to get highly reduced, so much that fields appear frozen. This goes on until the potential forces become non-negligible: this happens at some point during the radiation - matter phase. Then the scalar fields get accelerated and their kinetic energy starts rising again. It remains very small until after matter domination, when finally Hubble friction becomes small enough to allow a proper field displacement. Then, the fields roll down the potential, while dark energy is getting dominant. What happens exactly today and in the future is model dependent, but the whole dynamics just described for the past is (almost) not.
@@
Turning to the radiation - matter phase, we use the analytic expressions to obtain in (4.35) a field variation $\Delta\varphi$ during this period, that is
@@
This analytic expression, when evaluated on the examples, gives a very good match with the numerics. It shows in particular how frozen the fields are, namely $\Delta\varphi\lesssim 10^{-2}$. We further use the analytical solutions to get an expression (4.40) for $N_{{\rm kin}V}$, the moment at which $w_{\varphi}=0$, characterising its transition from $+1$ to $-1$: it is given by
@@
namely that the field distance in the matter - dark energy phase is sub-Planckian, as e.g. in Figure 2(a).
@@ 2.2 Equation system and reformulations
evolving in terms of $N$. The three equations can be used to form a system of first order differential equations. With a single canonical scalar field, an extra variable, $-\partial_{\varphi}V/V$, can be introduced. In the case where this variable is constant, it can be considered as a parameter. The system then becomes a closed first order differential system, that can be studied using the methods of dynamical systems. $-\partial_{\varphi}V/V$ being constant amounts to have an exponential potential; we refer to [15, (2.17)] for the corresponding system, and subsequent dynamical system analysis. Beyond the exponential potential however, this formulation is less interesting and we will rather use the above formulation, solving (2.20b), (2.20c).
@@ 3 Numerical solution examples, and domination phases
For $\Lambda$CDM, the solution is then completely fixed, but for the exponential and hilltop quintessence models, different solutions are obtained for different values of $w_{\varphi 0}$. In the following, we will fine-tune this value in order to get a solution with a radiation domination phase, that starts around $N=-20$. This is motivated by the time of BBN, which has the corresponding redshift $z_{{\rm BBN}}\sim e^{20}$ [15]; an earlier value for the start of the radiation phase is possible, by a further tuning of $w_{\varphi 0}$. Without such a fine-tuned value of $w_{\varphi 0}$, most of the solutions obtained do not have a radiation domination phase; those are thus unrealistic (see e.g. [15, Sec. 3.1]).
@@ 3.1 $\Lambda$CDM
From Figure 6, we learn that $H/H_{0}$ takes huge values in the early universe. This quantity decreases as expected (see Section 2.1), reaching in this model a value close to a constant today; the fact it asymptotes to a constant for $\Lambda$CDM can only be seen via the analytical solution, discussed in Section 4.4.1. Using some values of $N$ that will become relevant in Section 3.4, we get the following evaluation of the Hubble parameter
@@ 3.2 Exponential quintessence
We turn to the scalar field in Figure 8, not present for $\Lambda$CDM. We observe an initial phase in $\varphi^{\prime}(t)$ with a non-zero kinetic energy (here with $\varphi^{\prime}(t)<0$): [^3] The sign of $\varphi^{\prime}(t)$ in this very early universe is fixed here by the tuning of $w_{\varphi 0}$. A tiny change in that value can give a different solution with the opposite sign of $\varphi^{\prime}(t)$. This other solution is otherwise almost indistinguishable, since most quantities do not depend on this sign: see e.g. [15, Fig. 8, 17]. this corresponds to an initial kination phase that we will comment on in Section 3.4, and that can be seen already in Figure 7(a). We do not see this phase in $\varphi(t)$, due to numerical precision issues in the very initial moments. The freezing of the field, noticed in Figure 7(b), is barely visible in Figure 8(c): we notice a constant value of $\varphi(t)$ in the early universe, around the time when $\varphi^{\prime}(t)=0$.
@@ 4 Analytic solutions and properties
Having presented in Section 2 the equations to solve, and in Section 3 several examples of numerical solutions and their features, we tackle in this section the question of getting analytical solutions and proving some of their properties. We will address this question by focusing separately on the different phases of the universe history, listed in (3.17).
@@ 4.1 Single component solutions
To start with, several well-known analytical solutions (to $f_{1}=f_{2}=e^{i}=0$) are obtained by considering a single fluid component in the universe. We will see that they correspond to limits of the complete solutions within the corresponding component domination phase. In the dynamical system study of the exponential quintessence [15], these single component solutions correspond to fixed points; by extension here, we denote them as $P_{n}$ for the $n$-component. The solutions are the following
@@
In realistic cosmologies, $P_{r}$ and $P_{m}$ should be good approximations of complete solutions around their respective domination: we see for instance in Table 4 that $\Omega_{n}\approx 1$ and $\Omega_{n^{\prime}\neq n}\ll 1$ close to these maxima. [^4] Since complete quintessence solutions evolve away from these dominations before and after the maxima, we deduce that as fixed points, $P_{r}$ and $P_{m}$ cannot be stable but only saddle points, and cosmological solutions pass only close by. The latter is consistent with the dynamical system analysis of [15]. Similarly, $P_{{\rm kin}}^{\pm...\pm}$ should be the approximation of the complete quintessence solutions at early times, when kinetic energy dominates ($\Omega_{{\rm kin}}\approx 1$), and $P_{\Lambda}$ should play the same role for the asymptotic future in $\Lambda$CDM. We will verify these claims in the following when getting more complete analytical solutions.
@@ 4.2.2 Is kination necessarily happening?
The analytical solution found during kination - radiation domination required to neglect $V,\rho_{m},k$, compared to $\rho_{{\rm kin}},\rho_{r}$. The two densities $\rho_{m},\rho_{k}$ are naturally neglected given their scaling in $a$, that gets subdominant during an early kination - radiation phase. Having $V$ negligible during this phase is the condition that we will discuss below. In [15, Fig. 7], one can find examples where even in absence of a radiation domination phase (due to different initial conditions), kination can still take place. The argument given at the beginning of this subsubsection holds indeed independently of radiation.
@@
where the early times will correspond here to a large, negative $\varphi(t)$, since the early times speed is positive. We believe the converse statement is true but do not prove it here.
@@
where in the very early times limit, $\varphi(t)\sim-\infty$. In addition, the potential being negligible compared to kinetic energy at these moments gets rewritten as
@@
This concludes the proof on this bound on the potential at early times. ∎
@@
A consequence of the (to be proven) converse implication is that any bounded potential (e.g. a $\cos\varphi$) or even a power law, $V\sim\varphi^{n}$, allows for initial kination; in particular, this happens in our above examples of exponential or hilltop potentials. On the contrary, a consequence of the lemma (via the contrapositive implication) is that exponentials of exponentials, as in LVS scenarios of string compactifications [28], cannot provide an initial kination. We also note that $e^{-\lambda\varphi}$ with $\lambda\geq\sqrt{6}$ does not allow for a cosmological solution with past radiation domination together with today’s acceleration [15]. The latter being necessary to a realistic cosmology, an initial kination therefore appears likely in any realistic quintessence model and associated cosmological solutions.

### v2

@@ Abstract
We focus on minimally coupled (multi)field quintessence models, of thawing type, and their realistic solutions. In a model-independent manner, we describe analytically these cosmological solutions throughout the universe history. Starting with a kination - radiation domination phase, we obtain an upper bound on the scalar potential to guarantee an early kination: $V(\varphi)\ll e^{-\sqrt{6}\varphi}$. Turning to the radiation - matter phase, we obtain analytic expressions for the scale factor $a(t)$ (not $t(a)$) and the scalar fields $\varphi^{i}(t)$ (usually neglected). These allow us to evaluate analytically the freezing of scalar fields, typically $\Delta\varphi\lesssim 10^{-2}$, as well as the transition moment of the dark energy equation of state parameter $w_{\varphi}$ from $+1$ to $-1$, with excellent agreement to the numerics. We comment on this freezing in view of string theory model building, and of some cosmological events. Turning to the latest phase of matter - dark energy domination, we show that the (multi)field displacement is sub-Planckian: $\Delta\varphi\leq 1$. We also provide for that phase analytic expressions for $\int(w_{\varphi}+1)\,{\rm d}N$ in terms of matter evolution; we relate those to observational targets that we propose. Using finally the CPL parametrisation, while discussing a phantom behaviour, we derive analytic bounds on $w_{0}$ and $w_{a}$.
@@ 1 Introduction and results summary
For about a century, we have known that our universe is expanding. Less than thirty years ago, the present expansion was observed to be accelerating. This acceleration is due to a so-called dark energy, which behaves very differently than ordinary matter or radiation. What is the nature of dark energy? This open problem is a key question of modern fundamental physics. A standard proposal is to describe dark energy as a constant energy density given by a cosmological constant $\Lambda>0$. This has led to a concordance cosmological model named $\Lambda$CDM, well-tested and in fairly good agreement with observations. Another option however is to have an energy density varying with time, giving a dynamical dark energy. One realisation of this possibility is through so-called quintessence models [1, 2, 3, 4] (see e.g. [5, 6, 7, 8] for early reviews and constraints), where the dynamics is that of scalar fields evolving in a scalar potential. Recent (and coming) observations, as well as theoretical arguments, have revived the possibility of explaining dark energy by quintessence. In this work, we perform a model-independent analysis of realistic quintessence scenarios, identifying interesting theoretical features for model building, as well as characteristic signatures that could possibly be observed, helping to distinguish such models from $\Lambda$CDM.
On the observational side, the first results of the DESI collaboration have recently been released [9]. Those are based on collected data until redshift $z=4$ with an unprecedent precision level. Consistently with the recent DES results [10], the DESI observational results are compatible with a dynamical dark energy model. More DESI results are announced for the coming years, and further experiments (Euclid, LSST/Vera Rubin) will provide independent measurements to similar precision. On the theoretical side, difficulties have been pointed-out in obtaining a universe with a positive cosmological constant (a.k.a. a de Sitter spacetime) from string theory, a quantum gravity theory candidate to be a fundamental theory of Nature. These difficulties have been characterised in terms of the strong de Sitter conjecture of the swampland program [11, 12]. While not claimed to be impossible, there is no known well-controlled string theory construction of such a solution up-to-date. As an alternative, there has been a recent interest in rather realising quintessence models from string theory (see [13, 14, 15, 16, 17, 18, 19, 20, 21] for a recent sample, and [17] for further references). While necessary ingredients are naturally present in string effective theories, finding the right dynamics and an appropriate scalar potential is not an easy task. But a lot of activity is currently dedicated to these matters, including cosmological model building and related investigations [22, 23, 24, 25, 26, 27, 28].
@@
indicating a dominant dark energy component, that provides acceleration. We consider in Section 3 three different models, and for each of them we obtain a realistic solution. The first one is $\Lambda$CDM, which can be viewed as a particular case of (1.1), and the other two are single field quintessence models, with respectively an exponential or a hilltop scalar potential. While these two single field models are presented for illustration, most of our analytic results will be derived in multifield situations. For quintessence models, obtaining a realistic solution (especially with a radiation domination phase) requires a tuning of the initial conditions, which we achieve. The evolution of the $\Omega_{n}$ is then similar from one model to the other, and we display it in Figure 1. We indicate in Figure 1(a) that different phases in this universe history can be delimited by moments, expressed in terms of $N$, that characterise equality ($N_{{\rm kin}r},N_{rm},N_{m\varphi}$) or maxima ($N_{r},N_{m}$) of the $\Omega_{n}$. We discuss these phases and specific moments in great details in Section 3.4, and provide various related analytic expressions. Given some values today $\Omega_{n0}$, it can be seen in Figure 1(b) that some of these moments are necessarily different between $\Lambda$CDM and quintessence, leading us to introduce extra notations: for instance, we label the moment for the maximum of $\Omega_{m}$ as $N_{m\,\Lambda}$ or $N_{m\,q}$ respectively. Similarly for the equality moment, we distinguish $N_{m\varphi\,\Lambda}$ and $N_{m\varphi\,q}$.
@@
The evolution, depicted in Figure 2(b), indicates the initial domination of kinetic energy ($w_{\varphi}=1$) as in kination, followed by a domination of the scalar potential ($w_{\varphi}=-1$) as for a cosmological constant, and a final rise towards the value today. This recent evolution is the prime target of DESI observations and will be of interest here; in addition the transition moment at which $w_{\varphi}=0$ will be determined analytically. Note that the latest part of this evolution is that of thawing quintessence models, which are included in this study, and differs from that of freezing or tracker models [6].
@@
The analytic solutions obtained provide a better understanding of the physics of the scalar fields; that of the scale factor is nothing but a continuous growth at various rates. Provided some initial kinetic energy and a negligible potential, the fields start by rolling up or down the potential: this is kination. The Hubble friction being huge in the early universe, the fields are quickly slowed down. During radiation domination, the kinetic energy continues to get highly reduced, so much that fields appear frozen. This goes on until the potential forces become non-negligible: this happens at some point during the radiation - matter phase. Then the scalar fields get accelerated and their kinetic energy starts rising again. It remains very small until after matter domination, when finally Hubble friction becomes small enough to allow a proper field displacement. Then, the fields roll down the potential, while dark energy is getting dominant. What happens exactly today and in the future is model dependent, but the whole dynamics just described for the past is (almost) not.
@@
Turning to the radiation - matter phase, we use the analytic expressions to obtain in (4.35) - (4.36) the field variations during this period in a multifield setting, boiling down in the single field case to
@@
This analytic expression, when evaluated on the examples, gives a very good match with the numerics. It shows in particular how frozen the fields are, namely $\Delta\varphi\lesssim 10^{-2}$. We further use the analytic solutions to get an expression (4.42) for $N_{{\rm kin}V}$, the moment at which $w_{\varphi}=0$, characterising its transition from $+1$ to $-1$: it is given by
@@
namely that the field distance in the matter - dark energy phase is sub-Planckian, as e.g. in Figure 2(a). This is important in order to avoid possible quantum gravity corrections to the effective theory used, as argued e.g. through the refined swampland distance conjecture (see e.g. [29]).
@@ 2.2 Equation system and reformulations
evolving in terms of $N$. The three equations can be used to form a system of first order differential equations. With a single canonical scalar field, an extra variable, $-\partial_{\varphi}V/V$, can be introduced. In the case where this variable is constant, it can be considered as a parameter. The system then becomes a closed first order differential system, that can be studied using the methods of dynamical systems. $-\partial_{\varphi}V/V$ being constant amounts to have an exponential potential; we refer to [17, (2.17)] for the corresponding system, and subsequent dynamical system analysis. Beyond the exponential potential however, this formulation is less interesting and we will rather use the above formulation, solving (2.20b), (2.20c).
@@ 3 Numerical solution examples, and domination phases
For $\Lambda$CDM, the solution is then completely fixed, but for the exponential and hilltop quintessence models, different solutions are obtained for different values of $w_{\varphi 0}$. In the following, we will fine-tune this value in order to get a solution with a radiation domination phase, that starts around $N=-20$. This is motivated by the time of BBN, which has the corresponding redshift $z_{{\rm BBN}}\sim e^{20}$ [17]; an earlier value for the start of the radiation phase is possible, by a further tuning of $w_{\varphi 0}$. Without such a fine-tuned value of $w_{\varphi 0}$, most of the solutions obtained do not have a radiation domination phase; those are thus unrealistic (see e.g. [17, Sec. 3.1]).
@@ 3.1 $\Lambda$CDM
From Figure 6, we learn that $H/H_{0}$ takes huge values in the early universe. This quantity decreases as expected (see Section 2.1), reaching in this model a value close to a constant today; the fact it asymptotes to a constant for $\Lambda$CDM can only be seen via the analytic solution, discussed in Section 4.4.1. Using some values of $N$ that will become relevant in Section 3.4, we get the following evaluation of the Hubble parameter
@@ 3.2 Exponential quintessence
We turn to the scalar field in Figure 8, not present for $\Lambda$CDM. We observe an initial phase in $\varphi^{\prime}(t)$ with a non-zero kinetic energy (here with $\varphi^{\prime}(t)<0$): [^3] The sign of $\varphi^{\prime}(t)$ in this very early universe is fixed here by the tuning of $w_{\varphi 0}$. A tiny change in that value can give a different solution with the opposite sign of $\varphi^{\prime}(t)$. This other solution is otherwise almost indistinguishable, since most quantities do not depend on this sign: see e.g. [17, Fig. 8, 17]. this corresponds to an initial kination phase that we will comment on in Section 3.4, and that can be seen already in Figure 7(a). We do not see this phase in $\varphi(t)$, due to numerical precision issues in the very initial moments. The freezing of the field, noticed in Figure 7(b), is barely visible in Figure 8(c): we notice a constant value of $\varphi(t)$ in the early universe, around the time when $\varphi^{\prime}(t)=0$.
@@ 4 Analytic solutions and properties
Having presented in Section 2 the equations to solve, and in Section 3 several examples of numerical solutions and their features, we tackle in this section the question of getting analytic solutions and proving some of their properties. We will address this question by focusing separately on the different phases of the universe history, listed in (3.17).
@@ 4.1 Single component solutions
To start with, several well-known analytic solutions (to $f_{1}=f_{2}=e^{i}=0$) are obtained by considering a single fluid component in the universe. We will see that they correspond to limits of the complete solutions within the corresponding component domination phase. In the dynamical system study of the exponential quintessence [17], these single component solutions correspond to fixed points; by extension here, we denote them as $P_{n}$ for the $n$-component. The solutions are the following
@@
In realistic cosmologies, $P_{r}$ and $P_{m}$ should be good approximations of complete solutions around their respective domination: we see for instance in Table 4 that $\Omega_{n}\approx 1$ and $\Omega_{n^{\prime}\neq n}\ll 1$ close to these maxima. [^4] Since complete quintessence solutions evolve away from these dominations before and after the maxima, we deduce that as fixed points, $P_{r}$ and $P_{m}$ cannot be stable but only saddle points, and cosmological solutions pass only close by. The latter is consistent with the dynamical system analysis of [17]. Similarly, $P_{{\rm kin}}^{\pm...\pm}$ should be the approximation of the complete quintessence solutions at early times, when kinetic energy dominates ($\Omega_{{\rm kin}}\approx 1$), and $P_{\Lambda}$ should play the same role for the asymptotic future in $\Lambda$CDM. We will verify these claims in the following when getting more complete analytic solutions.
@@ 4.2.2 Is kination necessarily happening?
The analytic solution found during kination - radiation domination required to neglect $V,\rho_{m},k$, compared to $\rho_{{\rm kin}},\rho_{r}$. The two densities $\rho_{m},\rho_{k}$ are naturally neglected given their scaling in $a$, that gets subdominant during an early kination - radiation phase. Having $V$ negligible during this phase is the condition that we will discuss below. In [17, Fig. 7], one can find examples where even in absence of a radiation domination phase (due to different initial conditions), kination can still take place. The argument given at the beginning of this subsubsection holds indeed independently of radiation.
@@
where the early times will correspond here to a large, negative $\varphi(t)$, since the early times speed is positive. We believe the converse statement is true but do not prove it here. As side remarks, the early times refer here to the initial moments (where in the limit, quantities become infinite), so the lemma does not apply to a transient, finite duration kination phase. The idea here is simply that by definition, kinetic energy dominates the potential in these early moments, and we give the potential bound beyond which this cannot happen anymore. We also recall that such an early time kination is happening in our realistic quintessence solutions (see Figure 12).
@@
where in the very early times limit, $\varphi(t)\sim-\infty$. Note that an additional constant in the field can then be neglected. In addition, the potential being negligible compared to kinetic energy at these moments gets rewritten as
@@
Note that any finite scale $V_{0}$ in the scalar potential can be absorbed by shifting the scalar field in the left-hand side by a constant proportional to $\ln V_{0}$. As mentioned, such an additional constant to the field can be neglected in these early times; one is then left with the bare exponential. This concludes the proof on this bound on the potential at early times. ∎
@@
A consequence of the (to be proven) converse implication is that any bounded potential (e.g. a $\cos\varphi$) or even a power law, $V\sim\varphi^{n}$, allows for initial kination; in particular, this happens in our above examples of exponential or hilltop potentials. On the contrary, a consequence of the lemma (via the contrapositive implication) is that very steep potentials such as exponentials of exponentials, appearing e.g. in certain regimes of LVS [32] or KKLT [33] scenarios of string compactifications, cannot provide an initial kination. We also note that $e^{-\lambda\varphi}$ with $\lambda\geq\sqrt{6}$ does not allow for a cosmological solution with past radiation domination together with today’s acceleration [17]. The latter being necessary to a realistic cosmology, an initial kination therefore appears likely in any realistic quintessence model and associated cosmological solutions.